OpenStudy (anonymous):

** Advance Algebra With Financial App ** Does anyone know the formula to this problem ? Bianca purchased a home entertainment center for $3,987 using an 18-month deferred payment plan with an interest rate of 28.12%. She did not make any payments during the deferment period. What will Bianca’s monthly payment be if she must pay off the home entertainment center within three years after the deferment period? $110.75 $168.04 $250.62 $257.91

OpenStudy (anonymous):

is it this A = P(1 + r)^t ?

OpenStudy (anonymous):

Is this right ? A=3,987(1+0.2812)^(18)

OpenStudy (cwrw238):

compound interest

OpenStudy (anonymous):

is my formula right ?

ganeshie8 (ganeshie8):

this looks like annuity to me

OpenStudy (anonymous):

@ganeshie8 I found the right formula but I dont know what is n A = P(1+r/n)^(n*t)

ganeshie8 (ganeshie8):

it wont work

ganeshie8 (ganeshie8):

lookup your notes for annuity formula

ganeshie8 (ganeshie8):

ive the formula, but im not getting answer thats in ur choices

OpenStudy (amistre64):

defered adds to the balance exponent, something like \[B_n=B_o(k)^{c+t}-P\frac{1-(k)^t}{1-k}\]

OpenStudy (anonymous):

im trying to solve but im getting frustrated

ganeshie8 (ganeshie8):

i got 1121.29 somehow

OpenStudy (anonymous):

@amistre64 B=? K=? C=? N=? T=? etc..

OpenStudy (amistre64):

\[B_n=B_o(k)^{c+t}-P\frac{1-(k)^t}{1-k}\] \[0=B_o(k)^{c+3*12}-P\frac{1-(k)^{3*12}}{1-k}\] \[0=B_o(k)^{1.5*12+3*12}-P\frac{1-(k)^{3*12}}{1-k}\] \[0=B_o(k)^{12(4.5)}-P\frac{1-(k)^{36}}{1-k}\] \[B_o(k)^{12(4.5)}=P\frac{1-(k)^{36}}{1-k}\]

OpenStudy (amistre64):

B is balance, k is the compounding setup 1+r/c t is generaly the years

OpenStudy (anonymous):

Is there an easier formula please ?

OpenStudy (amistre64):

maybe, but this is the one i developed from my own studies; the accounting course textbook liked to try to define a new formula for all sorts of situations, and i simply used the one

OpenStudy (amistre64):

the k is a clean up variable to make it simpler to read thru

OpenStudy (amistre64):

k = 1+.2812/12 = 1.0234 \[3987(k)^{54}=P\frac{1-(k)^{36}}{1-k}\]solve for P

OpenStudy (amistre64):

the formula is just: Remaining Balance = Balance Compounded - Annuity Payments

OpenStudy (amistre64):

adjust the balance compounded by the derfered period

OpenStudy (anonymous):

so k=1.0234

OpenStudy (amistre64):

approximately yes

OpenStudy (amistre64):

it appears that the answers want a more exacting form

OpenStudy (amistre64):

off by a few cents is all

OpenStudy (amistre64):

i just saw ganshes attempt, just forgot the compounding effect, .2812/12

ganeshie8 (ganeshie8):

yeah thanks for the catch amistre :) http://www.wolframalpha.com/input/?i=3987%281%2B.2812%2F12%29%5E36+%3D+x%28%281-%281%2B.2812%2F12%29%5E36%29%2F%281-%281%2B.2812%2F12%29%29%29 still its off by few dollars :\

OpenStudy (amistre64):

that 36 on the left side, needs to be adjust by 18 months is all, in that timeframe there was no payments being made

ganeshie8 (ganeshie8):

oh deferment period we cant ignore ?

OpenStudy (amistre64):

the beginning balance at the start of the payments is 3987(k)^(18); not 3987

OpenStudy (amistre64):

k^18 * k^36 = k^ 54

ganeshie8 (ganeshie8):

oh the interest accumulates during deferment period also is it

OpenStudy (amistre64):

yep :) this aint rooms to go lol

ganeshie8 (ganeshie8):

wow ! i get it the exact answer now :) i feel accomplished !! thanks amistre

OpenStudy (amistre64):

:) youre welcome

OpenStudy (anonymous):

@amistre64 Thank you soo much !

OpenStudy (amistre64):

youre welcome. i remember everyone in the class i took trying to remember all the different forumlas and such; i never worried about it :)

OpenStudy (anonymous):

so the answer is 168.04

OpenStudy (amistre64):

no, that is without the defered period; its more like 250